OpenStudy (anonymous):

Solve x2 = 14x – 49.

Parth (parthkohli):

Subtract both sides by 14x - 49. \(\Large \color{Black}{\Rightarrow x^2 - 14x + 49 = 0 }\) Solve with the quadratic formula.

Parth (parthkohli):

Quadratic formula is: \(\Large \color{Black}{\Rightarrow x = {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\)

OpenStudy (accessdenied):

You can also notice that it factors into a perfect square. (-7*-7=49, -7 + -7 = -14) (x - 7)^2 = 0 Which would lead to the same answer.

Parth (parthkohli):

Note that the formula applies to an equation in the form \(ax^2 + bx + c = 0\)

OpenStudy (anonymous):

use quadratic equation. =)